OpenStudy (jazzyluvsu123):

x=1-1/6 The answer to my problem is a fraction but I just keep getting .833 repeating. FAN AND MEDAL

OpenStudy (jaidenblow123):

wait is the fraction in the equation

OpenStudy (jazzyluvsu123):

yes 1/6

OpenStudy (p0sitr0n):

1-1/6=6/6-1/6=5/6

OpenStudy (jaidenblow123):

use reciprical

OpenStudy (p0sitr0n):

additionally, there is a way to convert a repeating decimal number into a fraction, but here it is easier to do it right away

OpenStudy (jaidenblow123):

the reciprical is the flipped fraction

OpenStudy (jazzyluvsu123):

tyty

OpenStudy (jaidenblow123):

im taking 8th grade math lol